See Table 2.5 LOADING... showing financial statement data and stock price data for Mydeco Corp. a. What were​ Mydeco's gross margins each​ year? b. Compare​ Mydeco's gross​ margin, EBIT​ margin, and net profit margin in and . Which margins​ improved?

See Table 2.5 showing financial statement data and stock price data for Mydeco Corp.


a. What were Mydeco's gross margins each year?


b. Compare Mydeco's gross margin, EBIT margin, and net profit margin in

and . Which margins improved?


table 2.5
Answer #1

20210605_145108-min.jpg


20210605_145118-min.jpg




A) gross margin: gross profit/ revenue (find in table 2.5)


2015: 212.1/ 397.7= 0.5333165703 round to 53.33%


2016: 191.6/ 366.1= 0.5233542748 round to 52.34%


2017: 220.5/ 422.7= 0.5216465578 round to 52.16%


2018: 268.2/ 515.4= 0.5203725262 rounded to 52.04%


2019: 312.1/ 608.9= 0.5125636393 rounded to 51.26%


B) EBIT margin= EBIT/ revenue


2015: 58.9/ 397.7= 0.1481015841 rounded to 14.81 %


2019: 69.3/ 608.9= 0.1138117918 rounded to 11.38 %


C) net profit margin: net profit (net income)/ revenue


2015: 16.7/ 397.7= 0.0419914508 rounded to 4.20 %


2019: 20.6/ 608.9= 0.0338314994 rounded to 3.38 %


none of the margins improved from 2015- 2019.

answered by: Andrew San Andres